Changes between Version 9 and Version 10 of FilamentSupply_RevA_Testing_2_7_20


Ignore:
Timestamp:
Feb 11, 2020, 1:35:14 PM (3 years ago)
Author:
Michael
Comment:

--

Legend:

Unmodified
Added
Removed
Modified
  • FilamentSupply_RevA_Testing_2_7_20

    v9 v10  
    1313[[Image(http://gauss.bu.edu/svn/tof-system/Gallery/2020-2-7/HVRise2.jpg, 15%, link=[http://gauss.bu.edu/svn/tof-system/Gallery/2020-2-7/HVRise2.jpg])]]
    1414[[BR]]These results are not exactly as expected but they may be explainable. For the HV Pulser the ON signal is used as part of the control for the output waveform. In the Firmware used for this test, the ON signal was either fixed HIGH, fixed LOW, or alternating HIGH/LOW. In the expected usage of the HV Pulser this signal goes HIGH before the NEG signal, stays HIGH for some time, goes LOW, then the POS signal goes LOW for one clock and the cycle repeats. In the next round of testing a switch will be used to modify the ON signal to the desired behavior and hopefully that will fix the rise and fall times of this signal. The fluctuation of the Vin coming from the HV Pulser is also assumed to be present because of the absence of a clamping circuit present in the original design of the HV Pulser. For the purposes of these tests the overshoot of Vin doesn't matter so this can be ignored for now. It is also possible that the HV Pulser and Filament Supply are interacting in a negative way but for now I'm assuming these behaviors can be fixed by modifying the Firmware used for testing and eventually adding a clamping circuit to the HV Pulser.
     15
    1516[[BR]]On the bright-side. The bottom of the Load still followed the Vin at a fixed distance which is the desired behavior of the Filament Supply
    1617